Four Balls of wool will make 8 knitted caps. How many balls of wool will Malcolm need if he wants to make 6 caps
1
Expert's answer
2012-12-18T10:31:59-0500
Let's denote B the number of knitted caps that can be produced from one wool ball. Then

4B = 8 ==> B = 8/4 = 2.

So, we can make two caps from one ball. Therefore, Malcolm needs

N = 6/B = 6/2 = 3

balls to make 6 caps.